# Quantum Bayes Classifiers and Their Application in Image Classification

### v1

@@ V.1 The Accuracy of Binary Classification
It can be seen from Fig. 9 that the symmetric-QBC shows a better classification performance in most cases. While as in shown in Fig. 10, the SPODE-QBC shows better results in the Fashion-MNIST dataset in most cases, and the symmetric-QBC also exhibits better classification results on some data.
@@
  3. 3.
    Some features extracted from local sampling and binarization procedure can not distinguish data from two classes very well. Other sampling and binarization methods could be considered.
@@ VI Discussions & Conclusions
In this paper, we study the constructions of QBCs using Bayesian networks. Based on four kinds of Bayesian networks, four QBCs are implemented, that is, the naïve-QBC, the SPODE-QBC with the feature in an image center as the ’superparent’, the TAN-QBC, and the symmetric-QBC.
We apply these QBCs to image classification and propose the image classification algorithm. To reduce the computational complexity, a local feature sampling method is designed to extract a few key features from a huge number of image pixels. By keeping key features, this method maintains high classification accuracies while reducing the number of features. Adjusting the sampling block size and utilizing different convolution kernels can further improve the accuracies of QBCs.
The classification effects of QBCs are verified on the MindQuantum platform for the MNIST and Fashion-MNIST datasets, respectively. The results show that QBCs perform well for image classification. We also compare QBCs with the classical Bayesian classifier and QCNN that use all features directly, which shows that QBCs using a few features still have advantages in some cases.
Unlike QCNN classifiers HurKim-435 ; CongChoi-181 , QBCs do not require a time-consuming training process. They only require to calculate the corresponding parameters statistically and load them into quantum circuits, while the classification decision can be made by measuring quantum circuits, which is faster, lower in computational complexity, and less resource-consuming than QCNN.
At present, there does not exist a general QBC that performs well with all types of data. Building QBCs based on alternative varieties of Bayesian networks that work well with various datasets would be intriguing.

### v2

@@ V.1 The Accuracy of Binary Classification
It can be seen from Fig. 10 that the symmetric-QBC shows a better classification performance in most cases. While as is shown in Fig. 11, the SPODE-QBC shows better results in the Fashion-MNIST dataset in most cases, and the symmetric-QBC also exhibits better classification results on some data.
@@
  3. (3)
    Some attributes extracted from the local sampling and binarization procedure can not distinguish data from two classes very well. Other sampling and binarization methods could be considered.
@@ VI Discussions & Conclusions
In this paper, we study the constructions of QBCs using Bayesian networks. Based on four kinds of Bayesian networks, four QBCs are implemented, that is, the naïve-QBC, the SPODE-QBC with the attribute in an image center as the ‘superparent’, the TAN-QBC, and the symmetric-QBC. We apply these QBCs to image classification and propose the image classification algorithm. To reduce the computational complexity, a local feature sampling method is designed to extract a few key attributes from a huge number of image pixels. By retaining essential attributes, this method achieves high classification accuracies while reducing the feature size. Adjusting the sampling block size and utilizing different convolution kernels can further improve the accuracies of QBCs.
The classification effects of QBCs are verified on the MindQuantum platform for the MNIST and Fashion-MNIST datasets, respectively. The results show that QBCs perform well for image classification. We also compare QBCs with the classical Bayesian classifier and QCNN that use much larger size of features, which shows that QBCs using a fewer size of features still have advantages in some cases. Unlike QCNN classifiers HurKim-435 ; CongChoi-181 , QBCs do not require a time-consuming training process. They only require to calculate the corresponding parameters statistically and load them into quantum circuits, while the classification decision can be made by measuring quantum circuits, which is faster, lower in computational complexity, and less resource-consuming than QCNN. At present, there does not exist a general QBC model that performs well with all types of data. Building QBCs based on alternative varieties of Bayesian networks that work well with various datasets would be intriguing.